OpenStudy (anonymous):

Solve -7(k+9)=9(k-5)-14k in steps?

OpenStudy (anonymous):

Distribute: \[-7k-63=9k-45-14k\] Simplify \[-7k-63=-5k-45\] \[-18=2k\] Solve: \[k=\frac{-18}{2}=-9\] Check: \[-7(-9+9)=9(-9-5)-14(-9)\] \[0=0\]

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

No problem.